### v1

# Exploring the Adversarial Vulnerabilities of Vision-Language-Action Models in Robotics

Abstract
Recently in robotics, Vision-Language-Action (VLA) models have emerged as a transformative approach, enabling robots to execute complex tasks by integrating visual and linguistic inputs within an end-to-end learning framework. While VLA models offer significant capabilities, they also introduce new attack surfaces, making them vulnerable to adversarial attacks. With these vulnerabilities largely unexplored, this paper systematically quantifies the robustness of VLA-based robotic systems. Recognizing the unique demands of robotic execution, our attack objectives target the inherent spatial and functional characteristics of robotic systems. In particular, we introduce an untargeted position-aware attack objective that leverages spatial foundations to destabilize robotic actions, and a targeted attack objective that manipulates the robotic trajectory. Additionally, we design an adversarial patch generation approach that places a small, colorful patch within the camera’s view, effectively executing the attack in both digital and physical environments. Our evaluation reveals a marked degradation in task success rates, with up to a 100% reduction across a suite of simulated robotic tasks, highlighting critical security gaps in current VLA architectures. By unveiling these vulnerabilities and proposing actionable evaluation metrics, this work advances both the understanding and enhancement of safety for VLA-based robotic systems, underscoring the necessity for developing robust defense strategies prior to physical-world deployments.
1 Introduction
“First directive: A robot cannot harm a human or, through inaction, allow a human to be harmed.”
— Finch [77]
Figure 1: Fatal failure rates resulting from different attacking schemes (i.e., UADA, UPA, and TMA) underscore vulnerabilities in VLA-based robotic systems.
In the movie Finch, set in a post-apocalyptic future, an intelligent robot navigates complex themes, underscoring the importance of interactive safety with its human master. Once speculative, this portrayal is now increasingly plausible with the rise of visually-conditioned language models (VLMs) [72, 81] capable of seamlessly interpreting both visual and linguistic contexts. A notable realization of this potential can be seen in Vision-Language Action (VLA) models [75, 16, 35, 76], which integrate VLMs into robotic systems to enable end-to-end learning that encompasses both high-level planning and the generation of low-level robot control actions. While VLA models demonstrate a promising step toward generalist robotic intelligence, they also introduce potential vulnerabilities that remain underexplored. This work addresses the pressing need to investigate the new attack surfaces associated with VLA-based robotic systems.
To gain deeper insights into these vulnerabilities, we analyze VLA-based models and general robotic operations, highlighting two key characteristics crucial for designing effective adversarial attacks. First, creating attack objectives for robotic systems requires consideration of the physical dynamics and constraints intrinsic to robotic movement. Conventional adversarial attacks, in this case, often fail to produce significant deviations in intended actions because they disregard these constraints. Second, VLA models generate control signals that function as a time series of K-class predictions (see §3.1), making it essential to design attacks that exploit the temporal dependencies within these sequences to cause substantial disruptions in robotic behavior. Achieving attacks that are effective in both digital simulations and real-world environments remains critical yet challenging.
To address these challenges, our work intensifies the adversarial threats posed to VLA-based systems by developing specialized attack objectives and designing effective attack methods. Specifically, we formulate an Action Discrepancy Objective aimed at maximizing the normalized action discrepancy within VLA-based robotic systems. This strategy ensures that, at each decision point, the robot’s behavior can diverge from the optimal trajectory. Additionally, we introduce Geometry-Aware Objective that considers the robot’s movement in three-dimensional space, characterized by three degrees of freedom. By optimizing the cosine similarity between adversarial and ground-truth directions, we induce deviations in the robot’s movement direction from its intended path, increasing the likelihood of task failure. To achieve these attack objectives, we develop a straightforward yet effective Patch-Based Attack targeting VLA-based robotic systems. This approach enables adversarial attacks in both digital and physical settings, revealing substantial vulnerabilities within the VLA-based system.
Together, these innovations yield several significant contributions: This work presents a pioneering and comprehensive analysis of vulnerabilities within VLA-based robotic systems, a new paradigm for training generalized robotic policies using generative foundation models. We reveal critical threats to adversarial attacks (see Fig. 1), emphasizing the urgent need to enhance robustness before real-world deployment. We are the first to define attack objectives specific to the powerful VLA model and to employ a straightforward adversarial patch against it (see §3). This offers valuable insights for the research community to explore systemic failures in similar concurrent generative foundation models. We rigorously evaluate our approach in both simulated and physical environments across four distinct robotic tasks, observing a significant raises in task failure rates of 100% and 43%, respectively. This highlights the effectiveness of our attack strategies (see §4.2).
Figure 2: Overall Adversarial Framework. The robot captures an input image, processes it through a vision-language model to generate tokens representing actions, and then uses an action de-tokenizer for discrete bin prediction. The model is optimized with adversarial objectives focusing on various discrepancies and geometries (i.e., UADA, UPA, TMA). Forward propagation is shown in black, and backpropagation is highlighted in pink. These objectives aim to maximize errors and minimize task performance, with visual emphasis on 3D-space manipulation and a focus on generating adversarial perturbation $\delta$ during task execution, such as picking up a can.
2 Related work
2.1 AI-driven Generalist Robot
Developing generalist robots [62, 6, 38, 27, 14] requires models not only handle varied interactions but also maintain robustness in unstructured environments. Early intelligent robots [56, 52, 53, 55, 85] relied on rule-based approaches, effective only in controlled settings and requiring extensive reprogramming for new tasks. The advent of deep learning [41, 58, 19, 31, 33] and reinforcement learning [65, 74, 1] has shifted this paradigm to adapt through data, increasing robot versatility and reducing manual reconfiguration.
Recent studies [7, 54, 59] have further advanced the development and potential of large-scale visually-conditioned language models (VLMs) as key enablers for generalist robots, demonstrating promising generalization across a variety of scenes [86, 32, 66, 25, 42, 15]. One of the notable examples is OpenVLA [35], which integrates dual visual encoders with a pre-trained large language model to enable observation-to-action capabilities through instruction-following [46, 47, 34, 10, 84, 72]. By aligning visual and textual semantics, OpenVLA facilitates complex scenario understanding and end-to-end action generation, showing impressive generalization when executing previously unseen task instructions. Despite its success in generalist robotic applications, the vulnerability of such systems remains largely unexplored. This study aims to address this gap by investigating adversarial threats to VLA-based robotic systems.
2.2 Human-Robot Interaction Safety
Current human-robot interaction risks include physical discomfort or injury [22, 4, 5]. Considering these risks, ensuring safety during interactions becomes the top priority as many literature outlined [69, 24, 40]. Existing robotic safety strategies have primarily concentrated on mitigating naturally occurring hazards [40]. These strategies can be generally categorized into pre-collision or post-collision interventions based on the timing of manage hazards [39]. Pre-collision approaches aim to prevent severe hazards by implementing physical constraints, such as controlling force and speed [28, 30], defining safety zones [70] and predicting human actions to avoid potentially dangerous robot movements [37, 36, 51]. Post-collision strategies, on the other hand, focus on detection and response after a harmful action has occurred [49, 20]. While there is substantial research on robotic safety, the integration of AI into robotic policies poses new challenges, where attackers can manipulate AI-driven robot and execute malicious actions. Such threats introduce additional layers of risk and complexity. In this paper, we primarily focus on adversarial threats.
2.3 Adversarial Attacks in Robotics
Adversarial attacks serve as critical tools for assessing model vulnerabilities, particularly in robotics, where models operate in dynamic, real-world settings. Traditional gradient-based, pixel-level attacks [21, 50, 17, 73, 44, 63] exploit model gradients to calculate malicious perturbations, achieving high attack success rates in digital environments. For physical-world attacks, patch-based methods [8, 61, 79, 3, 78] offer a practical alternative for real-world applications by introducing physically realizable perturbations that retain efficacy under varied conditions, such as angle or lighting changes, making them highly applicable for robotic systems. Therefore, considering VLA-based models enable robot execution in real-world scenarios, this work employs a patch-based attack strategy.
VLA models [66, 26, 13] couple visual and linguistic modalities for perception-action alignment. The continuous, high-dimensional nature of visual inputs [64, 50, 60] makes them the most likely target for adversarial perturbations, as attackers can subtly alter visual data in ways that are difficult to detect [9, 2, 67]. Consequently, this work applies attacks on the visual modality to achieve robust manipulations across diverse environments. To the best of our knowledge, this study is one of the earliest attempts to formally define adversarial objectives for VLA models, targeting gaps in adversarial robustness through the model’s geometric constraints (e.g., spatial dependencies in visual input) and architectural nuances (e.g., cross-modal information integration).
3 Methodology
This section outlines our proposed methodology. We first review the core algorithmic principles of VLA in §3.1, which serves as the foundation for the adversarial scheme designs. We then present two types of untargeted adversarial attacks, including Untargeted Action Discrepancy Attack (UADA) in §3.2 and Untargeted Position-aware Attack (UPA) in §3.3. In §3.4, we further include Targeted Manipulation Attack (TMA), designed to direct the robot toward specific erroneous execution. The implementation details are presented in §3.5. The overall framework is shown in Fig. 2.
3.1 Preliminary
Vision-Language Action models [86, 35] (see Fig. 2) are built on large language models integrated with visual encoders, enabling robots to interpret human instructions and process visual input from a camera to perform context-aware actions. VLA models [35, 86, 43] generally abstract continuous action prediction into a classification problem by discretizing robot actions. Specifically, they first discretize the continuous probabilities into class labels $c=\arg\max F(x)$. An action de-tokenizer then generates actions $y=DT(c)$. By categorizing action values into discrete class labels, the model converts continuous probability outputs into discrete signals, this simplification facilitates quicker convergence and faster training times, and is commonly used for VLA-based models [35, 86, 43].
In this work, we focus on a 7-degree-of-freedom (DoF) robotic arm [23]. At each step, an action consists of 7DoFs with specific physical significance in three-dimensional Cartesian space, represented by:
$\small A=[\Delta{P_{x}},\Delta{P_{y}},\Delta{P_{z}},\Delta{R_{x}},\Delta{R_{y}},\Delta{R_{z}},\text{gripper}],$ (1)
where $\Delta{P_{x,y,z}}$ and $\Delta{R_{x,y,z}}$ denote relative positional and rotational changes along the x-, y-, and z-axis, discretized into 256 uniform bins across the action range [35]. The $gripper$ state is binary, indicating its open or closed state. This control design presents a unique challenge for adversarial attacks, as finely divided bins result in minimal action discrepancies between neighboring bins (e.g., $\pm$0.007/bin). This means that even if an attack shifts the classification to an adjacent bin, the resulting action discrepancy has minimal impact real-world performance.
3.2 Untargeted Action Discrepancy Attack
We first focus on maximizing action discrepancies by introducing Untargeted Action Discrepancy Attack (UADA). This attack is based on the observation that larger robot actions usually correlate with intense physical movements, which, in turn, may amplify the potential for real-world hazards [28, 30, 29]. Specifically to UADA, the attack target is one or combination of 7DoFs, defind as $y^{I}=\{y^{i}|i\in[1,\dots,7]\}$. To define UADA’s objective, we first determine the maximum action discrepancy $d_{\text{max}}$, which represents the largest deviation that can occur within the allowed range for each DoF. This is calculated by measuring the distance between the ground truth $y$ and its action range $[y_{\min},y_{\max}]$ as:
$\small d_{max}^{i}(y)=\max(|y^{i}-y^{i}_{\min}|,|y^{i}-y^{i}_{\max}|).$ (2)
We next calculate the applied action discrepancy $d_{\text{applied}}^{i}(x,y)$ $=|F(x)^{i}-y^{i}|$ in order to measure the deviation between the model’s output and the ground truth, where $F(x)^{i}$ is the $i$-th action output of model $F(x)$. In this sense, we can evaluate the Normalized Action Discrepancy (NAD) as the drifting ratio between $d_{applied}^{i}$ and $d_{max}^{i}$ as:
$\rm{NAD}=\frac{1}{I}\sum_{i}^{I}d^{i}_{\text{applied}}/d^{i}_{\text{max}}.$ (3)
Since VLA models are currently operated via predicting discrete token labels (see §3.1), we then need to bridge action discrepancy derived from these discrete labels into learnable targets. Specifically, we weighted sum the model output probabilities across $J$ bins by their own class label $j$ as:
$d_{\text{applied}}^{i\prime}(x,y)=|\sum_{j=1}^{J}(j\cdot\mathcal{F}(x)_{j}^{i})-c^{i}|,$ (4)
where $\mathcal{F}(\cdot)_{j}^{i}$ is the model output probability of $i$-th DoF’s $j$-th bin. Finally, the attack objective for UADA is defined as minimizing the ratio of weighted sum $d^{\prime}_{\text{applied}}$ to $d_{max}(y^{i})$:
$\small\mathcal{L}_{\text{UADA}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}\left[\sum^{I}_{i}\frac{d^{i\prime}_{\text{applied}}(x+\delta,y)}{d^{i}_{\max}(y)}\right]^{-1},$ (5)
where $x,y$ is a benign image and action pair, sampled from distribution $\mathcal{X}$. UADA defines the objective functions tailored for the robot’s action space, allowing adversarial patches to create significant, lasting disruptions in task performance. In this design, we uniquely leverage the model’s discrete action structure to design attacks that amplify meaningful discrepancies in robot behavior.
3.3 Untargeted Position-aware Attack
We further extend untargeted adversarial attacks to consider the positional dynamics within VLA models. In typical task execution, precise and directed movements of the end-effector towards designated goals are essential, with actions mapped in three-dimensional space. The positional degrees of freedom, $y^{pos}=[y^{1},y^{2},y^{3}]$ encapsulates the directional movements within 3D for effective robotic control.
Recognizing the importance of $y^{pos}$ in controlling the end-effector’s path, we introduce a position-aware attack to disrupt the intended movement trajectory. This type of adversarial vulnerability remains largely unexplored but has significant implications for task failure: the attack objective can steer the end-effector away from its intended path by introducing directional perturbations, amplifying errors and causing substantial trajectory distortions. ${}_{\!}$we define the Untargeted Position-aware Attack (UPA) objective as:
$\small\mathcal{L}_{\text{UPA}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}\big{[}\alpha\frac{y^{pos}_{adv}\cdot y^{pos}}{\|y^{pos}_{adv}\|\|y^{pos}\|}+\beta\frac{1}{||y^{pos}_{adv}-y^{pos}||_{2}}\big{]},$ (6)
where $\alpha$ and $\beta$ are the hyperparameters that balance between the directionality and magnitude of the perturbations. By integrating geometric awareness into our attack, this approach generates perturbations that induce cumulative deviations in the robot’s trajectory, even with minimal adjustments. $L_{2}$-normalization $||\cdot||_{2}$ further intensifies these deviations, rendering the attack highly effective and disruptive.
1:  Input: $\mathcal{X}$: dataset; $\delta$: adversarial patch; $\mathcal{L}_{o}$: objective function; $\mathcal{F}$: vision language action model; $\mathcal{T}(\cdot)$: Transfromation; $\phi,\psi$ are two hyperparameters for transformation; T: maximum iteration steps; $k$: maximum inner-loop steps.
2:  Initialize: $\delta\sim\mathcal{U}[0,1)$, $\mathcal{L}_{o}\in\{\mathcal{L}_{\text{UADA}},\mathcal{L}_{\text{UPA}},\mathcal{L}_{\text{TMA}}\}$
3:  for $i=1,2,\cdots,T$ do
4:     $(x,y)\sim\mathcal{X}$
5:     for $i=1,2,\cdots,k$ do
6:        $shx,shy\sim\mathcal{U}(-\phi,\phi)$, $\theta\sim\mathcal{U}(-\psi,\psi)$
[…]
10:        Clip $\delta$ into normal pixel range.
11:     end for
12:  end for
Algorithm 1 Adversarial Patch Attack.
Figure 3: Qualitative Results of adversarial vulnerabilities over OpenVLA-7B [35] and OpenVLA-7B-LIBERO [35] with objectives of UADA, UPA, and TMA, respectively. We visualize the overall 3D trajectories and 2D trajectories of benign $\bullet$ and adversarial $\bullet$ scenarios at each time step to compare the impact of the generated adversarial patch in affecting them. The untargeted trajectory $\bullet$ is visualized in UADA task. All trajectories start with , and we plot the success end point, marked as .
3.4 Targeted Manipulation Attack
In addition to the aforementioned untargeted attacks, we also explore the adversarial vulnerabilities of VLA models with a targeted strategy. This attack aims to mislead the models into predicting specific trajectories, resulting in precise alterations. These targeted alterations can potentially drive malicious behaviors or induce task failure. The objective of the targeted manipulation attack is:
$\small\mathcal{L_{\text{TMA}}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}[\mathcal{H}(\mathcal{F}(x+\delta),y^{I}_{T})],$ (7)
where $y^{I}_{T}=\{y^{i}=t|i\in[1,\dots,7],t\in\mathcal{R}\}$ represents the adversarial target; and $\mathcal{H}(\cdot,\cdot)$ denotes the Cross-Entropy loss, which measures the error between the predicted state under adversarial perturbation and the desired target state. By steering the robot towards an adversarial target across successive time steps, our approach manipulates the trajectory and undermines task performance, ultimately leading to a substantial disruption of intended operations. Furthermore, manipulating any individual DoF or a combination thereof can also severely compromise task success, given that the errors in one dimension tend to propagate and magnify over time during execution (see Fig. 3).
3.5 Implementation Details
The implementation of our pipeline is detailed in Algorithm 1. We incorporate two key modifications to enhance the training stability of the generated patch.
Inner-loops. Following previous work [50], we add $k$ inner-loops optimize steps during each iteration in line 5, aiming to reduce data variance and ensure more consistent updates.
Geometric Transformations. To improve the robustness of our attack under real-world scenarios, we employ random geometric transformations $\mathcal{T}(\cdot)$ in line 1 with affine and rotation transformations [3] as:
$\small\mathcal{T}(\cdot)=\begin{bmatrix}1&sh_{y}\\ sh_{x}&1\end{bmatrix}\cdot\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix},$ (8)
where $sh_{x},sh_{y}\sim\mathcal{U}(-\phi,\phi)$ are shear factors, and $\theta\sim\mathcal{U}(-\psi,\psi)$ denotes the rotation angle, both sampled from uniform distributions during inner loop optimization step $k$.
4 Experiments
In this section, we present the experimental design outlined in §4.1 and report the quantitative and qualitative results of UADA, UPA, and TMA in both simulation and physical-world scenarios (§4.2). We then analyze diagnostic experiments (§4.3) to evaluate the influence of key components. Additionally, we assess the robustness of our method against various defense mechanisms in §4.4 and conclude with a comprehensive systemic discussion in §4.5.
4.1 Experiment Setup
Simulation Tasks. We conduct simulation attacks on BridgeData V2 [71] and LIBERO [45]. BridgeData V2 [71] is a real-world dataset comprising $24$ diverse environments and $13$ distinct skills, such as grasping, placing, and object rearrangement, with a total of $60,096$ trajectories. LIBERO [45] is a simulation dataset designed to evaluate models across four distinct task types (Spatial, Object, Goal, Long), each containing $10$ tasks with $50$ human-teleoperated demonstrations. Notably, LIBERO-Long combines diverse objects, layouts, and long-horizon tasks, making them challenging for complex, multi-step planning, thus we choose LIBERO-Long to conduct UADA, UPA and TMA.
Robot Setups. For the physical-world tasks, we adopt a robotic system consisting of a Universal Robots UR10e equipped with a Robotiq Hand-E Gripper to provide a 7DoF motion. The sensing system includes one RGB webcam in the fixed position with a shoulder view.
Victim VLA Models. In our study, we select publicly available and current most influential VLAs as victim models for evaluation. Specifically, we focus on four variants of the OpenVLA model, each trained independently on different tasks within the LIBERO dataset (i.e., Spatial, Object, Goal, Long). To evaluate the effectiveness of our methods, we craft adversarial patches using two distinct generating setups. The first involves a model trained in a simulated environment using the LIBERO dataset [45] with the openvla-7B-libero-long variant [35]. The second uses a model trained on physical world data from the BridgeData v2 dataset [71] with the openvla-7B model [35]. This approach allows us to assess our methods on both simulated and real-world data. Subsequently, we evaluate the performance of generated adversarial patches on victim models (i.e., OpenVLA LIBERO variants) trained on different tasks to rigorously prove the robustness and effectiveness of our method. These models are trained with distinct data sources and task objectives. Additional results on transferability (e.g., LLaRA [43]) are included in §S1.
Attacking Details. Aligned with the former objective [68] aiming to mislead classifiers and combine with our framework, we design and conduct an untargeted attack. Following [35], we apply scaling and normalization for data preprocessing. The initial learning rate is set to $2e^{-3}$, with the AdamW optimizer [48] with a cosine annealing scheduler, including $20$ warm-up iterations. Training is conducted over $2e^{3}$ iterations with a batch size of $6$ and $50$ inner-loop steps (see Algorithm 1). The transformation parameters $a,b$ in Eq. 8 are set to $0.2$ and $30$ to simulate real-world perspective change. For all three objectives, considering the trade-off between stealth and performance observed in §S2, we set the patch size to $5\%$. For the hyperparameters $\alpha,\beta$ in Eq. 6, We conduct grid search and tune $\alpha$ from $[0.2,0.4,0.6,0.8]$ and $\beta$ from $[0.8,0.6,0.4,0.2]$. We select $\alpha=0.8$ and $\beta=0.2$ based on the validation set performance. Ablation studies $w.r.t.$ $\alpha$ and $\beta$ are presented in §S6.
Table 1: Untargeted results (Simulation). We report Failure rates (%) and NAD (%) in LIBERO simulation. denotes the in-domain victim model and dataset aligned with the patch generation model and dataset, denotes a transfer attack task involving a distinct victim model and dataset (see §S5). The naive FR of each victim model in the corresponding task is: Spatial (15.3%), Object (11.6%), Goal (20.8%) and Long (46.3%). The FR ($\uparrow$) is highlighted in best and second best. Same for Tab. 2.
Victim models Objective Attack Actions NAD (%) Spatial Object Goal Long Untarget DoF1 14.1 100.0 100.0 100.0 100.0     DoF1-3 11.4 100.0 100.0 100.0 100.0 UADA DoF1 18.1 100.0 100.0 100.0 100.0     DoF1-3 14.3 100.0 100.0 100.0 100.0 UPA     DoF1-3 14.5 100.0 83.4 93.4 96.7
Table 2: Untargeted results (Physical). Transfer results to 4 victim models with adversarial patches generated from physical data and corresponding model.
Victim model Objective Attack Actions NAD (%) Spatial Object Goal Long Untarget DoF1 26.7 85.4 56.4 77.3 65.8     DoF1-3 23.8 78.1 46.7 52.4 56.7 UADA DoF1 31.9 73.4 33.5 41.7 62.3     DoF1-3 26.4 85.9 49.5 32.1 60.5 UPA     DoF1-3 26.9 86.4 42.7 53.3 61.2
Evaluation Details. To assess the effectiveness and robustness of our methods, we conduct evaluations on the LIBERO dataset [45]. Each task suite is evaluated with $10$ trials, with fixed patch location (see §S5) This setup is for avoiding the patch obstructing task-related objects.
Building on the concept of robotic tasks Success Rate (SR) introduced in LIBERO [45], we adopt Failure Rate (FR), defined as $1-\text{SR}$, as the primary evaluation metric for this study. To further quantify action discrepancy, we employ NAD (see Eq. 3) for untargeted attacks, as it considers relative distances to quantify deviation from the optimal action. A higher NAD value indicates a greater discrepancy between the ground truth and the predicted action. For targeted attacks, we utilize the L1 distance, which measures the absolute deviation from the specified target action.
Reproducibility. Our methods are implemented in Pytorch [57]. Experiments are conducted on a single NVIDIA A100-40GB GPU. Our code and attack video is available at https://anonymous.4open.science/r/roboticAttack-56FF.
Table 3: Targeted Manipulation Attack results. Average failure rate (%) are reported in 4 LIBERO simulation tasks. The performance of AFR ($\uparrow$) is marked in best and second best. The initial AFR across the four tasks is 23.5%.
Generation Setup Metric Target Actions Simulation / DoF1 DoF2 DoF3 DoF4 L1 0.022 0.029 0.040 0.029 AFR 97.8 97.5 96.8 73.4 / … DoF5 DoF6 DoF7 DoF1-3 L1 0.024 0.023 0.046 0.031 AFR 96.8 88.3 84.1 95.5 Physical / DoF1 DoF2 DoF3 DoF4 L1 0.013 0.022 0.041 0.031 AFR 89.1 85.3 57.2 48.8 / … DoF5 DoF6 DoF7 DoF1-3 L1 0.032 0.018 0.049 0.027 AFR 54.3 46.2 57.8 86.8
4.2 Main Result
Quantitative Results. For UADA and UPA, our methods effectively amplify action discrepancies, leading to a notable transfer attack ability in increasing failure rates (see Tab. 1). Specifically, for the Simulation setup, UADA and UPA achieve maximum NAD of 18.1% and 14.5%, significantly outperforming untargeted scenarios with increment of 3.8% and 0.4%, respectively. Both untargeted and UADA successfully disrupt the robot execution, with a failure rate to 100%. UPA also results in a significant increment compared to the failure rate of the victim methods in benign scenarios, which is 100% $v.s.$ 15.3%, 83.4% $v.s.$ 11.6%, 93.4% $v.s.$ 20.8% and 96.7% $v.s.$ 46.3%, respectively. Regarding the Physical setup, both UADA and UPA generate actions with large action discrepancies, which are 31.9% and 26.9% with an increase of 5.2% and 0.2% compared to the best 26.7% in untargeted scenarios, respectively. UPA demonstrates an improvement of 86.4% $v.s.$ 85.9% over the spatial task compared to untargeted.
For TMA task, we evaluate the effectiveness of our method by manipulating various DoF to 0. Our method demonstrates significant effectiveness in manipulating robotic trajectory and increasing failure rates (see Tab. 3). Specifically, when applied to the different generation settings, our approach yields notable increments across various tasks, including an max average failure rate 97.8% $v.s.$ 23.5% of benign performance in Simulation, and 89.1% in Physical, respectively. To further prove the generalizability, we target attack on several non-zero values in §S2.
Qualitative Results. We qualitatively analyze the trajectories of robot movement of the proposed three different objectives in Fig. 3. For UADA, we compare the trajectory deviations in the same trail with patch generated from UADA and untarget attack. As seen, the untarget attack induces small deviations in the trajectory UADA, on the other hand, produces significantly larger trajectory deviations (also supported by NAD metric in Tab. 2), which is attributed to UADA’s capability of incorporating action discrepancies. In our observation, UADA significantly amplifies the effect on overall task execution, thereby increasing the potential for robotic hazards. For UPA, we observe chaotic and irregular behaviors, including instances where the end-effector moves out of the camera’s field of view. We attribute this phenomenon to the efficacy of the adversarial patch in perturbing the model’s spatial perception, inducing a consistent deviation from the intended direction of movement, ultimately resulting in failure in a long run.
For TMA, we observe a notable reduction in the range of motion along the x-axis corresponding to the targeted attack axis. This suggests that our proposed targeted attack can effectively constrained robot movements.
In summary, the qualitative analysis shows that both Untarget attack, UADA, UPA, and TMA can effectively disrupt the robot actions generated from VLA models. These findings underscore a pressing security concern during the deployment of generalist robots, especially when considering application scenes that requires reliable operations [11, 70].
Figure 4: Impact of Inner-loop, Patch Size and Defense Discussion. The figure shows how varying Inner-loop affects NAD in UADA, and patch sizes affect $L1$ distance and the failure rates in TMA, both targeting at DoF1. (a) Impact of Inner-loop, (b) Impact of Patch Size and (c-f) the effect of four different defenses on failure rates.
Figure 5: Qualitative Results of the physical world. The first/second row show benign and adversarial cases respectively.
Physical-world Performance. In addition to the digital simulation results, we also conduct a comprehensive evaluation of the performance of our generated adversarial patches in physical-world testing scenarios. The evaluation encompassed 100 trials across three distinct tasks: object grasping, placement, and manipulation (see §S3 for more details). As shown in Fig. 5, the adversarial patch generated with UADA demonstrated the ability to manipulate the robot effectively, achieving an attack success rate exceeding 43%. Although this success rate is lower than the corresponding digital-world performance (i.e., 85.9%), it highlights the effectiveness of our patches in physical-world applications as well without the need for further adaptations. Crucially, we observed that the induced erratic movements (similar to simulation scenario) of the robotic manipulators during successful attacks pose significant risks to human safety and surrounding environments. This finding emphasizes the severe threat posed by VLA models in real-world applications, consistent with similar observations in digital settings.
4.3 Diagnostic Experiment
We ablate components of our method in this section.
Impact of Inner-loops. We discuss the impact of inner-loop steps to model performance in Fig. 4(a). The results show that NAD first improves when inner-loop steps continue to increase. This is due to the reduced data variance and gradually stabilized gradients. However, further increasing the steps results in a slower performance increase, with a noticeably longer training schedule. We thus choose inner-loop steps=50 as it balances between performance and scale.
Impact of Patch Size. We further study TMA’s performance with different patch sizes, and report the average failure rate across four LIBERO [45] benchmark tasks in Fig. 4. The patch sizes examined were $[3\%,5\%,7\%,10\%]$ of a $224\times 224$ input image. Our findings indicate an inversely proportional relationship between the $L_{1}$ distance of predicted actions and the target action as the patch size increased. This increase in patch size also correlated with a marked rise in task failure rates. The observed trend suggests that larger patch sizes afford the adversary more extensive optimization space to influence the model’s output, which is consistent with previous work [12, 83].
4.4 Robustness Evaluation
We further examine if concurrent defense strategies [18, 82, 80] can generalize against our adversarial examples introduced within VLA. Specifically, we applied four prior defense techniques, and reported their task failure rates for both benign and adversarial samples in Fig. 4, respectively. The results indicate that our adversarial attack can escape from most of the defense strategies. More details of defense parameters see in §S7.
4.5 Systemic Discussion
(a) BridgeData V2 [71]
(b) LIBERO [45]
Figure 6: Patch Visualization. Sematic-rich patches generated from physical or simulation datasets.
Patch Pattern Analysis. In Fig.6, we present the adversarial patches generated for a range of targets. A particularly noteworthy observation is that some of these patches bear a striking resemblance to the structural joints of a robotic arm. Given the strong similarity between these patches and robotic arms in appearance, we hypothesize that these VLA-based models, in their current training paradigms, are often limited to a single robotic system operating within restricted camera views. This constrained perspective/camera pose may inadvertently induce a learned representation bias, where adversarial perturbations align with prominent structural features of the training robot’s appearance in order to deceive current victim models. Given that these patches are model-specific and reflect the spatial and visual patterns of the training environment, we have a reasonable concern that they may compromise VLA-based models’ generalizability and robustness in physical-world deployments.
Possible Defense Strategies. Our work highlights the pressing need for a paradigm shift in the VLA-based models’ training strategies in order to successfully defend our attacks. Specifically, future approaches might incorporate more training scenarios that involve complex multi-robot interactions to mitigate the current patch bias. Another possible solution is to leverage the logical relationships within the robot arm’s physical structure to estimate the actual position of each joint. By predicting joint positions deviated from physical constraints, we can eliminate unreasonable adversarial samples, even if they are similar to actual joint(s).
5 Conclusion
While VLA models have gained significant popularity due to their substantial robot capabilities, this study, for the first time, establishes specific robotic attack objectives and demonstrates that these models are in fact vulnerable to various types of attacks. Experimental results demonstrate that our attacks expose significant vulnerabilities to VLA models, raising concerns regarding impetuous real-world deployments. We believe our framework provides pioneering and foundational contributions to the reliability of AI enhanced robotic systems.
6 Acknowledgements
This research was supported by the National Science Foundation under Grant No. 2242243. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of U.S. Naval Research Laboratory (NRL) or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.