### v3

# Exploring the Adversarial Vulnerabilities of Vision-Language-Action Models in Robotics

Abstract
Recently in robotics, Vision-Language-Action (VLA) models have emerged as a transformative approach, enabling robots to execute complex tasks by integrating visual and linguistic inputs within an end-to-end learning framework. While VLA models offer significant capabilities, they also introduce new attack surfaces, making them vulnerable to adversarial attacks. With these vulnerabilities largely unexplored, this paper systematically quantifies the robustness of VLA-based robotic systems. Recognizing the unique demands of robotic execution, our attack objectives target the inherent spatial and functional characteristics of robotic systems. In particular, we introduce two untargeted attack objectives that leverage spatial foundations to destabilize robotic actions, and a targeted attack objective that manipulates the robotic trajectory. Additionally, we design an adversarial patch generation approach that places a small, colorful patch within the camera’s view, effectively executing the attack in both digital and physical environments. Our evaluation reveals a marked degradation in task success rates, with up to a 100% reduction across a suite of simulated robotic tasks, highlighting critical security gaps in current VLA architectures. By unveiling these vulnerabilities and proposing actionable evaluation metrics, we advance both the understanding and enhancement of safety for VLA-based robotic systems, underscoring the necessity for continuously developing robust defense strategies prior to physical-world deployments [^1] Our code is available at https://github.com/William-wAng618/roboticAttack. Project Homepage at https://vlaattacker.github.io/.
1 Introduction
“First directive: A robot cannot harm a human or, through inaction, allow a human to be harmed.”
— Finch [75]
Figure 1: Adversarial Vulnerabilities induced by malicious manipulation. (A). Illustration of adversarial threats in robotic task execution. (B). Example of semantic-rich adversarial patches generated by proposed methods. (C). Comparison of failure rates across different attack schemes (UADA, UPA, and TMA).
In the movie Finch, set in a post-apocalyptic future, an intelligent robot navigates complex themes, underscoring the importance of interactive safety with its human master. Once speculative, this portrayal is now increasingly plausible with the rise of Large Vision Language Models (LVLMs) [70, 79], capable of seamlessly interpreting both visual and linguistic contexts. A notable realization of this potential can be seen in Vision-Language Action (VLA) models [73, 16, 34, 74], which integrate LVLMs into robotic systems to enable end-to-end learning that encompasses both high-level trajectory planning and low-level robot action control. While VLA models demonstrate a promising step toward generalist robotic intelligence, they also introduce potential vulnerabilities that remain largely underexplored. This work, in light of this view, pioneers the understandings of the pressing need to investigate the new attack surfaces associated with VLA-based robotic systems.
To gain deeper insights into these vulnerabilities, we analyze VLA-based models and general robotic operations, highlighting two key characteristics crucial for designing effective adversarial attacks. First, creating attack objectives for robotic systems requires consideration of the physical dynamics and constraints intrinsic to robotic movement. Conventional adversarial attacks, in this case, often fail to produce significant deviations in intended actions because they disregard these constraints. Second, VLA models generate control signals that function as a time series of K-class predictions (see §3.1), making it essential to design attacks that exploit the temporal dependencies within these sequences to cause substantial disruptions in robotic behavior. Achieving attacks that are effective in both digital simulations and real-world environments remains critical yet challenging.
To address these challenges, our work intensifies the adversarial threats posed to VLA-based systems by both developing specialized attack objectives and designing effective attack methods. Specifically, we formulate an Action Discrepancy Objective aimed at maximizing the action discrepancy within VLA-based robotic systems. This strategy ensures that, at each decision point, the robot’s behavior can diverge from the optimal trajectory. Additionally, we introduce Geometry-Aware Objective that considers the robot’s movement in three-dimensional space, characterized by three degrees of freedom. By optimizing the cosine similarity between adversarial and ground-truth directions, we induce deviations in the robot’s movement direction from its intended path, increasing the likelihood of task failure. To achieve these attack objectives, we develop straightforward yet effective Patch-Based Attacks targeting VLA-based robotic systems. This approach enables adversarial attacks in both digital and physical settings, revealing substantial vulnerabilities within the VLA-based system.
Altogether, these innovations yield several significant contributions: This work presents a pioneering and comprehensive analysis of vulnerabilities within VLA-based robotic systems, a new paradigm for training generalized robotic policies using generative foundation models. We reveal critical threats to adversarial attacks, emphasizing the urgent need to enhance robustness before real-world deployment; To the best of our knowledge, we are the first work to define attack objectives specific to the powerful VLA models and to employ a straightforward adversarial patch against it (see §3). This offers valuable insights for the research community to explore systemic failures in similar concurrent generative foundation models; We rigorously evaluate our approach in both simulated and real-world environments across four distinct robotic tasks, observing a significant raises in task failure rates of 100% and 43%, respectively. This highlights the effectiveness of our attack strategies (see §4.3).
Figure 2: Overall Adversarial Framework. The robot captures an input image, processes it through a vision-language model to generate tokens representing actions, and then uses an action de-tokenizer for discrete bin prediction. The model is optimized with adversarial objectives focusing on various discrepancies and geometries (i.e., UADA, UPA, TMA). Forward propagation is shown in black, and backpropagation is highlighted in pink. These objectives aim to maximize errors and minimize task performance, with visual emphasis on 3D-space manipulation and a focus on generating adversarial perturbation $\delta$ during task execution, such as picking up a can.
2 Related work
2.1 AI-driven Generalist Robot
Developing generalist robots [60, 6, 37, 27, 14] requires models to not only handle varied interactions but also maintain robustness in unstructured environments. Early intelligent robots [55, 51, 52, 54, 83] generally relied on rule-based approaches, effective only in controlled settings and needing extensive reprogramming for new tasks. The advent of deep learning [40, 56, 19, 31] and reinforcement learning [63, 72, 1] has shifted this paradigm to adapt through the data-driven pipeline, increasing robot versatility and reducing manual reconfiguration.
Recent studies [7, 53, 57] have further advanced the development and potential of Large Vision Language Models (LVLMs) as key enablers for generalist robots, demonstrating promising generalization across a variety of scenes [84, 32, 64, 25, 41, 15]. One of the notable examples is OpenVLA [34], which integrates dual visual encoders with a pre-trained large language model to enable observation-to-action capabilities through instruction-following [45, 46, 33, 10, 82, 70]. By aligning visual and textual semantics, OpenVLA facilitates complex scenario understanding and end-to-end action generation, showing impressive generalization when executing previously unseen task instructions.
2.2 Adversarial Attacks in Robot
Adversarial attacks serve as critical tools for assessing model vulnerabilities, particularly in robotics, where models operate in dynamic, real-world settings. Traditional gradient-based, pixel-level attacks [21, 49, 17, 71, 43, 61] exploit model gradients to calculate malicious perturbations, achieving high attack success rates in digital environments. For physical-world attacks, patch-based methods [8, 59, 77, 3, 76] offer a practical alternative for real-world applications by introducing physically realizable perturbations that retain efficacy under varied conditions, such as angle or lighting changes, making them highly applicable for robotic systems. Therefore, considering VLA-based models enable robot execution in real-world scenarios, this work employs a patch-based attack strategy.
VLA models [64, 26, 13] couple visual and linguistic modalities for perception-action alignment. The continuous, high-dimensional nature of visual inputs [62, 49, 58] makes them the most likely target for adversarial perturbations, as attackers can subtly alter visual data in ways that are difficult to detect [9, 2, 65]. Consequently, this work applies attacks on the visual modality to achieve robust manipulations across diverse environments. To the best of our knowledge, this study is one of the earliest attempts to formally define adversarial objectives for VLA models, targeting gaps in adversarial robustness through the model’s geometric constraints (e.g., spatial dependencies in visual input) and architectural nuances (e.g., cross-modal information integration).
3 Methodology
In this section, we first review the core algorithmic principles of VLA in §3.1, which serves as the foundation for the adversarial scheme designs. We then present two types of untargeted adversarial attacks, including Untargeted Action Discrepancy Attack (UADA) in §3.2 and Untargeted Position-aware Attack (UPA) in §3.3. In §3.4, we further include Targeted Manipulation Attack (TMA), designed to direct the robot toward specific erroneous execution. Finally, we introduce the Normalized Action Discrepancy (NAD) metric in §3.5. The overall framework is shown in Fig. 2.
3.1 Preliminary
Vision-Language Action models [84, 34] (see Fig. 2) are built on large language models integrated with visual encoders, enabling robots to interpret human instructions and process visual input from a camera to perform context-aware actions. VLA models [34, 84, 42] generally abstract continuous action predictions into a classification problem by discretizing robot actions. Specifically, they first discretize the continuous probabilities into class labels $y=\arg\max\mathcal{F}(x)$ where $\mathcal{F(\cdot)}$ is the VLA model. An action de-tokenizer then generates actions $A=DT(y)$. By categorizing action values into discrete class labels, the model converts continuous probability outputs into discrete signals, this simplification facilitates quicker convergence and faster training times, and is commonly used for VLA-based models [34, 84, 42].
In this work, we focus on a 7 degree-of-freedoms (DoFs) robotic arm [23]. At each step, an action consists of 7 DoFs with specific physical significance in three-dimensional Cartesian space, represented by:
$\small A=[\Delta{P_{x}},\Delta{P_{y}},\Delta{P_{z}},\Delta{R_{x}},\Delta{R_{y}},\Delta{R_{z}},\text{gripper}],$ (1)
where $\Delta{P_{x,y,z}}$ and $\Delta{R_{x,y,z}}$ denote relative positional and rotational changes along the x-, y-, and z-axis, discretized into 256 uniform bins across each DoFs’ action bounds $[y_{\min}^{i},y_{\max}^{i}]$ [34]. The $gripper$ state is binary, indicating its open or closed state. This control design presents a unique challenge for adversarial attacks, as finely divided bins result in minimal action discrepancies between neighboring bins (e.g., $\pm$0.007/bin). This means that even if an attack shifts the classification to an adjacent bin, the resulting action discrepancy has minimal impact on real-world performance.
3.2 Untargeted Action Discrepancy Attack
To exacerbate action discrepancies, we introduce the Untargeted Action Discrepancy Attack (UADA), which aims to maximize deviations in robot actions. This attack is based on the observation that larger robot actions usually correlate with intense physical movements, which, in turn, may amplify the potential for real-world hazards [28, 30, 29]. For UADA, the attack target is one or a combination of 7 DoFs, defined as $y^{I}_{gt}=\{y^{i}_{gt}|i\in[1,\dots,7]\}$. To define UADA’s objective, we first identify the most distant action $y_{adv}^{i}$, which maximizes the discrepancy from the $i$-th DoF ground truth action $y^{i}$. This action, for each DoF, corresponds to its own action bound $[y_{\min}^{i},y_{\max}^{i}]$ as:
$y_{adv}^{i}=\begin{cases}y_{\max}^{i},&\text{if }|y_{\max}^{i}-y^{i}_{gt}|\geq|y_{\min}^{i}-y^{i}_{gt}|\\ y_{\min}^{i},&\text{otherwise}\end{cases}.$ (2)
Instead of directly using $y_{adv}^{i}$ as the misclassification target, we introduce a soft attack objective to capture the discrepancy between actions, ensuring smooth gradient optimization and stable attack performance. Specifically, due to the physical action magnitude information contains in bin labels, we reweight the output probability $F(x)^{i}_{bins}\in\mathcal{R}^{{}_{1\times{bins}}}$ using normalized bin labels $y^{i}_{bins}=[\frac{1}{bins},\dots,1]$, where $\lfloor y^{i}_{bins}\rfloor$ and $\lceil y^{i}_{bins}\rceil$ correspond to the normalized values of $y_{\min}^{i}$ and $y_{\max}^{i}$, respectively. The reweighting process is defined as:
$y_{soft}^{i}=\sum_{bins=1}^{n}F(x+\delta)^{i}_{bins}\otimes y^{i}_{bins},$ (3)
where $\otimes$ denotes Hadamard Product, $\delta$ is the adversarial patch and $y_{soft}^{i}$ represents the $i$-th DoF’s soft action. Finally, the objective of UADA is to minimize the discrepancy between the soft and the most distant actions, which is:
$\mathcal{L}_{\text{UADA}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}\sum_{i}^{I}(y_{soft}^{i}-y_{adv}^{i})^{2},$ (4)
UADA defines the objective function $\mathcal{L}_{\text{UADA}}$ tailored for attacking the robot’s action space, allowing adversarial patches to create significant, lasting disruptions in task performance.
3.3 Untargeted Position-aware Attack
We explore untargeted adversarial attacks in parallel by considering the positional dynamics within VLA models. In typical task execution, precise and directed movements of the end-effector towards designated goals are essential, with actions mapped in three-dimensional space. The positional DoFs, $A^{p}=[\Delta{P_{x}},\Delta{P_{y}},\Delta{P_{z}}]$ encapsulates the directional movements within 3D for effective robotic control.
Recognizing the importance of $A^{p}=DT(y^{p})$ in controlling the end-effector’s path, we introduce a position-aware attack to disrupt the intended movement trajectory. This type of adversarial vulnerability remains largely unexplored but has significant implications for task failure: the attack objective can steer the end-effector away from its intended path by introducing directional perturbations, amplifying errors and causing substantial trajectory distortions. ${}_{\!}$we define the Untargeted Position-aware Attack (UPA) objective as:
$\small\mathcal{L}_{\text{UPA}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}\big{[}\alpha\frac{y^{p}_{adv}\cdot y^{p}}{\|y^{p}_{adv}\|\|y^{p}\|}+\beta\frac{1}{||y^{p}_{adv}-y^{p}||_{2}}\big{]},$ (5)
where $\alpha$ and $\beta$ are the hyperparameters that balance between the directionality and magnitude of the perturbations. By integrating geometric awareness into our attack, this approach generates perturbations that induce cumulative deviations in the robot’s trajectory, even with minimal adjustments. $L_{2}$-normalization $||\cdot||_{2}$ further intensifies these deviations, rendering the attack highly effective and disruptive.
Figure 3: Qualitative Results of adversarial vulnerabilities over OpenVLA-7B [34] and OpenVLA-7B-LIBERO [34] with objectives of UADA, UPA, and TMA, respectively. We visualize the overall 3D trajectories and 2D trajectories of benign $\bullet$ and adversarial $\bullet$ scenarios at each time step to compare the impact of the generated adversarial patch in affecting them. The untargeted trajectory $\bullet$ is visualized in UADA task. All trajectories start with , and we plot the success end point, marked as .
3.4 Targeted Manipulation Attack
In addition to the aforementioned untargeted attacks, we also explore the adversarial vulnerabilities of VLA models with a targeted strategy. This attack aims to directly mislead the models into predicting specific trajectories, resulting in precise alterations. We design this attack since the targeted alterations can drive malicious behaviors or induce task failure. The objective of the targeted manipulation attack is:
$\small\mathcal{L_{\text{TMA}}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}[CE(\mathcal{F}(x+\delta)^{I},y^{I}_{T})],$ (6)
where $y^{I}_{T}=\{y^{i}_{T}=t|i\in[1,\dots,7],t\in[y^{i}_{\min},y^{i}_{\max}]\}$ represents the adversarial target(s); and $CE(\cdot,\cdot)$ denotes the Cross-Entropy loss, which measures the error between the predicted state under adversarial perturbation and the desired target state. By steering the robot toward an adversarial target across successive time steps, our approach manipulates the trajectory and undermines task performance, ultimately leading to a substantial disruption of intended operations. Furthermore, manipulating any individual DoF or a combination thereof can also severely compromise task success, given that the errors in one dimension tend to propagate and magnify over time during execution (see Fig. 3).
3.5 A Novel Evaluation Metric:
Normalized Action Discrepancy
To assess the potential physical impact of adversarial patches, we introduce Normalized Action Discrepancy (NAD) as a fine-grained metric for evaluating discrepancies at the action level. NAD serves as a complementary measurement alongside the coarse-grained Failure Rate (FR) at the task level, providing a detailed analysis of deviations during action execution. To define the NAD, we first determine the maximum action discrepancy $d^{i}_{\text{max}}$, which represents the largest deviation that can occur within the allowed range for $i$-th DoF. This is computed by measuring the L1 distance between the ground truth $y^{i}_{gt}$ and its action bound $[y^{i}_{\min},y^{i}_{\max}]$ as:
$\small d_{\text{max}}^{i}(y)=\max\big{[}DT(|y^{i}_{gt}-y^{i}_{\min}|),DT(|y^{i}_{gt}-y^{i}_{\max}|)\big{]}.$ (7)
We next calculate the applied action discrepancy $d_{\text{applied}}^{i}(x,y)$ $=|DT(F(x)^{i})-DT(y^{i}_{gt})|$ to measure the deviation between the model’s output and ground truth. We define the NAD as the drifting ratio between $d_{\text{applied}}^{i}$ and $d_{\text{max}}^{i}$ as:
$\rm{NAD}=\frac{1}{I}\sum_{i}^{I}\frac{d^{i}_{\text{applied}}}{d^{i}_{\text{max}}},$ (8)
where $I\in[1,\dots,7]$ represents a single or a combination of DoF(s). In this way, NAD provides normalized measurement of action deviations, allowing for a consistent evaluation of adversarial impact across varying DoF(s).
4 Experiments
In this section, we first detail the implementation of our adversarial framework and baseline methods in §4.1. Next, we outline the experimental setup in §4.2 and present quantitative and qualitative results in both simulation and physical-world scenarios (§4.3). We then conduct diagnostic experiments (§4.4) to analyze the impact of key components. Additionally, we assess the robustness of our method against various defense mechanisms in §4.5 and conclude with a comprehensive systemic discussion in §4.6.
1:  Input: $\mathcal{X}$: dataset; $\delta$: patch; $\mathcal{L}_{o}$: attack objective; $\mathcal{F}$: VLA model; $\mathcal{T}(\cdot)$: transformations; $\phi,\psi$: transforma -tion params; T: attack steps; $k$: inner-loop steps.
2:  Initialize: $\delta\sim\mathcal{U}[0,1)$, $\mathcal{L}_{o}\in\{\mathcal{L}_{\text{UADA}},\mathcal{L}_{\text{UPA}},\mathcal{L}_{\text{TMA}}\}$
3:  for $i=1,2,\cdots,T$ do
4:     for $i=1,2,\cdots,k$ do
5:        $sh_{x},sh_{y}\sim\mathcal{U}(-\phi,\phi)$, $\theta\sim\mathcal{U}(-\psi,\psi)$
6:        $y_{\text{pred}}\leftarrow\mathcal{F}\big{[}\mathcal{T}((x+\delta),(sh_{x},sh_{y},\theta))\big{]}$
[…]
8:     end for
9:  end for
Algorithm 1 Adversarial Patch Attack.
4.1 Implementation Details
The implementation of our adversarial patch attack pipeline is detailed in Algorithm 1. We incorporate two key modifications to enhance the training stability of the generated patch.
Inner-loops. Following previous work [49], we add $k$ inner-loops optimize steps during each iteration in line 4, aiming to reduce data variance and ensure more consistent updates.
Geometric Transformations. To improve the robustness of our attack under real-world scenarios, we employ random geometric transformations $\mathcal{T}\big{[}(\cdot),(sh_{x},sh_{y},\theta)\big{]}$ in line 1 with affine and rotation transformations [3] as:
$\small\mathcal{T}\big{[}(\cdot),(sh_{x},sh_{y},\theta)\big{]}=\begin{bmatrix}1&sh_{y}\\ sh_{x}&1\end{bmatrix}\cdot\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix},$ (9)
where $sh_{x},sh_{y}\sim\mathcal{U}(-\phi,\phi)$ are shear factors, and $\theta\sim\mathcal{U}(-\psi,\psi)$ denotes the rotation angle, both sampled from uniform distributions during inner loop optimization step $k$.
Untargeted Manipulation Attack (UMA) Baseline. As a pioneering work, we could not find a directly comparable baseline for our attacks. Therefore, we adapt prior work in adversarial learning as one of our baseline methods [66]. This attack is designed to mislead classifiers; we integrate it into our framework, leading to the following objective:
$\small\mathcal{L_{\text{UMA}}}=-\mathbb{E}_{(x,y)\sim\mathcal{X}}[CE(\mathcal{F}(x+\delta)^{I},y^{I}_{gt})].$ (10)
By minimizing the negative CE loss, this attack pushes the predicted action away from the ground truth action.
Random Noise Baseline. We generate random noise patches as an additional baseline, representing an unstructured intervention without any learned adversarial intent. Specifically, the noise is sampled from a Gaussian distribution $\mathcal{N}(0,1)$ and evaluated with the same setting as proposed attacks.
4.2 Experiment Setup
Datasets. We conduct attacks on BridgeData V2 [69] and LIBERO [44] with corresponding VLAs. BridgeData V2 [69] is a real-world dataset comprising $24$ diverse environments and $13$ distinct skills, such as grasping, placing, and object rearrangement, with a total of $60,096$ trajectories. LIBERO [44] is a simulation dataset designed to evaluate models across four distinct task types (Spatial, Object, Goal, Long). Notably, LIBERO-Long combines diverse objects, layouts, and long-horizon tasks, making them challenging for complex, multi-step planning, thus we choose LIBERO-Long to conduct UADA, UPA and TMA.
Robot Setups. For the real-world tasks, we adopt a robotic system consisting of a Universal Robots UR10e equipped with a Robotiq Hand-E Gripper to provide a 7DoF motion. The sensing system includes one RGB webcam in the fixed position with a shoulder view.
Victim VLAs. In our study, we select publicly available and current most influential VLAs as victim models for evaluation. Specifically, we focus on four variants of the OpenVLA model, each trained independently on different task suites within the LIBERO dataset [44] (i.e., Spatial, Object, Goal, Long). To evaluate the effectiveness of our methods, we craft adversarial patches using two distinct generating setups: Simulation Setting involves a model trained in a simulated environment using the LIBERO Long suite with the openvla-7B-libero-long variant [34]. Physical Setting uses a model trained on physical world data from the BridgeData v2 [69] with the openvla-7B model [34]. This approach allows us to assess our methods on both simulated and real-world data. Subsequently, we evaluate the performance of generated adversarial patches on victim models (i.e., OpenVLA LIBERO variants) trained on different tasks suites to rigorously prove the robustness and effectiveness of our method. These models are trained with distinct data sources and task objectives. Additional results on transferability (e.g., LLaRA [42]) are included in §S2.
Table 1: Untargeted results (Simulation). We report Failure Rates (FR) and NAD in LIBERO simulation. denotes the in-domain victim model and dataset aligned with the patch generation model and dataset, denotes a transfer attack task involving a distinct victim model and dataset (see §4.2). The FR ($\uparrow$) is highlighted in best and second best for each task suite. Same for Tab. 2.
Victim Model Objective Action(s) NAD(%) Spatial Object Goal Long Avg Benign - - 15.3$\pm$10.211.6$\pm$10.020.8$\pm$12.046.3$\pm$18.623.5% Random Noise 28.8$\pm$24.214.8$\pm$7.921.0$\pm$15.548.4$\pm$14.828.3% UMA DoF1 14.1 100$\pm$0.099.0$\pm$3.0100$\pm$0.0100$\pm$0.099.8%     DoF1∼3 11.4 100$\pm$0.098.8$\pm$2.699.0$\pm$2.4100$\pm$0.099.5% UADA DoF1 21.0 100$\pm$0.099.2$\pm$2.4100$\pm$0.0100$\pm$0.099.8%     DoF1∼3 17.9 100$\pm$0.0100$\pm$0.0100$\pm$0.0100$\pm$0.0100% UPA     DoF1∼3 14.5 96.2$\pm$11.477.8$\pm$12.588.0$\pm$10.496.8$\pm$3.089.7%
Evaluation Details. To assess the effectiveness and robustness of our methods, we conduct experiments on the LIBERO dataset [44]. Each suite consists of 10 tasks, with each task executed for 50 trials, resulting in a total of 500 rollouts, following Kim et al. [34]. For reproducibility, we carefully select patch paste locations for each suite, ensuring they do not obscure objects or robots in the test environment. Further details are provided in §S3.
Evaluation Metric. Regarding the task execution evaluation, we take the maximum steps of each task suite in the LIBERO training dataset as the timeout failure condition to reduce computational overhead. Furthermore, building on the concept of Success Rate (SR) introduced in LIBERO [44], we adopt Failure Rate (FR), defined as $\small 1-\text{SR}$, as the primary evaluation metric. To further quantify action discrepancy, we employ NAD (see Eq. 8) for untargeted attacks on corresponding DoF(s), as it considers relative distances to quantify deviation from the optimal action. A higher NAD value indicates a greater discrepancy between the ground truth and the predicted action. Additionally, we report the standard deviation of FR across tasks within each task suite.
Table 2: Untargeted results (Physical). Transfer results to 4 victim models with adversarial patches generated from physical settings. We report Failure Rates (FR) and NAD. FR ($\uparrow$) and NAD ($\uparrow$) are highlighted in best and second best for each task suite.
Victim Model Objective Action(s) NAD(%) Spatial Object Goal Long Avg Benign - - 15.3$\pm$10.211.6$\pm$10.020.8$\pm$12.046.3$\pm$18.623.5% Random Noise 28.8$\pm$24.214.8$\pm$7.921.0$\pm$15.548.4$\pm$14.828.3% UMA DoF1 26.7 89.6$\pm$15.360.8$\pm$16.864.8$\pm$26.880.0$\pm$20.873.8%     DoF1∼3 23.8 96.6$\pm$6.256.4$\pm$20.480.0$\pm$20.582.0$\pm$17.978.8% UADA DoF1 32.6 99.2$\pm$1.898.8$\pm$2.492.6$\pm$15.296.6$\pm$4.896.8%     DoF1∼3 27.7 100$\pm$0.0100$\pm$0.0100$\pm$0.0100$\pm$0.0100% UPA     DoF1∼3 26.9 95.6$\pm$9.5 % 57.0$\pm$23.657.2$\pm$24.472.8$\pm$19.770.7%
Table 3: Targeted Manipulation Attack results. Failure Rate (FR, $\uparrow$) and its standard deviation across tasks within LIBERO [44] suite are reported. The performance of task suits and DoFs is averaged separately, with the best and second-best values separately.
Simulation (Attacked on OpenVLA-LIBERO-10) Physical (Attacked on OpenVLA-7b) Targeted DoF(s) Metric Spatial Object Goal Long Avg Spatial Object Goal Long Avg Benign FR 15.3$\pm$10.211.6$\pm$10.020.8$\pm$12.046.3$\pm$18.623.5% 15.3$\pm$10.211.6$\pm$10.020.8$\pm$12.046.3$\pm$18.623.5% Random Noise 28.8$\pm$24.214.8$\pm$7.921.0$\pm$15.548.4$\pm$14.828.3% 28.8$\pm$24.214.8$\pm$7.921.0$\pm$15.548.4$\pm$14.828.3% DoF1 FR 100$\pm$0.097.8$\pm$3.0100$\pm$0.0100$\pm$0.099.5% 90.6$\pm$19.936.6$\pm$18.161.0$\pm$21.786.4$\pm$12.768.7% DoF2 100$\pm$0.099.2$\pm$1.3100$\pm$0.0100$\pm$0.099.8% 99.6$\pm$0.869.2$\pm$11.176.2$\pm$26.191.8$\pm$9.984.2% DoF3 100$\pm$0.097.2$\pm$6.0100$\pm$0.0100$\pm$0.099.3% 72.0$\pm$14.830.2$\pm$17.439.2$\pm$20.962.4$\pm$17.551.0% DoF4 94.0$\pm$11.037.4$\pm$23.251.2$\pm$26.587.6$\pm$14.967.6% 93.6$\pm$9.731.8$\pm$20.345.0$\pm$18.363.8$\pm$17.058.6% DoF5 100$\pm$0.092.6$\pm$8.4100$\pm$0.0100$\pm$0.098.2% 97.8$\pm$4.936.0$\pm$22.562.0$\pm$24.879.4$\pm$21.368.8% DoF6 100$\pm$0.080.0$\pm$9.592.6$\pm$13.4100$\pm$0.093.2% 79.0$\pm$14.222.2$\pm$14.441.0$\pm$21.562.6$\pm$15.751.2% DoF7 99.2$\pm$1.386.6$\pm$8.367.2$\pm$22.596.4$\pm$6.487.4% 91.8$\pm$15.398.8$\pm$1.370.6$\pm$23.493.2$\pm$10.288.6%  DoF1∼3 100$\pm$0.0100$\pm$0.0100$\pm$0.0100$\pm$0.0100% 96.4$\pm$8.083.6$\pm$16.874.4$\pm$26.391.8$\pm$11.986.6% Task Avg 99.2% 86.4% 88.9% 98.0% - 90.1% 51.1% 58.7% 78.9% -
4.3 Main Result
Quantitative Results. For UADA and UPA, our methods effectively amplify action discrepancies, leading to a notable transfer attack ability in increasing failure rates (see Tab. 1). Specifically, while attacking DoF1 and DoF1∼3 in the Simulation setup, UADA and UPA achieve NAD of 21.0% and 14.5%, significantly outperforming UMA scenarios with increments of 6.9% and 3.1%, respectively. Both UADA and UPA effectively disrupt robot execution, yielding maximum average failure rates of 100% and 89.7%, respectively. Regarding the Physical setup (see Tab. 2), our attack methods demonstrate strong transferability, as adversarial patches generated in the physical setting significantly impact the performance of LIBERO-X test. Specifically, both UADA and UPA generate malicious actions with large action discrepancies, which are 32.6% (DoF1) and 26.9% (DoF1∼3) with an increase of 5.9% and 3.1% compared to 26.7% (DoF1) and 23.8% (DoF1∼3) in UMA scenarios, respectively. Compare the two settings (Tab. 1 $v.s.$ Tab. 2), we observe a general variance in NAD (e.g. 32.6% $v.s.$ 21.0% for UADA). This discrepancy can be attributed to the fundamental differences between the simulation and real-world datasets used for conducting attacks (Bridge V2 [69] $v.s.$ LIBERO [44]). The increased variability in real-world data, including environmental complexity, object diversity, and task difficulty, allows the robot more opportunities to generate larger action discrepancies within the validation dataset.
For TMA task, we evaluate the effectiveness of our method by manipulating all DoF(s) to 0 (i.e., $y^{i}_{T}=0$ in Eq. 6). Our method demonstrates significant effectiveness in manipulating robotic trajectory and increasing FR (see Tab. 3). Specifically, when applied to the different generation settings, our approach yields notable increments across various tasks, including a max average failure rate 100% $v.s.$ 23.5% of benign performance in Simulation, and 88.6% in Physical, respectively. Moreover, during attack DoF4, we observe low FR and high task deviation. This failure can be attributed to the fact that DoF4 controls the orientation along the x-axis, which can be redundant DoF in tasks. As a result, attacks targeting redundant DoF(s) [38] are less likely to disrupt execution effectively. This observation underscores the importance of task-specific considerations when designing adversarial attacks on robotic systems. To further prove the generalizability, we target on several non-zero values in §S4.
Figure 4: Qualitative Results of the physical world. The first/second row show benign and adversarial cases respectively.
Figure 5: Impact of Inner-loop, Patch Size and Defense Discussion. The figure shows how varying Inner-loop affects NAD in UADA, and patch sizes affect $L1$ distance and the failure rates in TMA, both targeting at DoF1. (a) Impact of Inner-loop, (b) Impact of Patch Size and (c-f) the effect of four different defenses on failure rates.
Qualitative Results. We qualitatively analyze robot movement trajectories under the three proposed attacks in Fig. 3. For UADA, we compare the trajectory deviations in the same trail with patches generated from UADA and UMA attacks. As seen, the UMA induces small deviations in the trajectory. UADA, on the other hand, produces significantly larger trajectory deviations (also supported by NAD metric in Tab. 2), which is attributed to UADA’s capability of incorporating action discrepancies. In our observation, UADA significantly amplifies the effect on overall task execution, thereby increasing the potential for robotic hazards. For UPA, we observe chaotic and irregular behaviors, including instances where the end-effector moves out of the camera’s field of view. We attribute this phenomenon to the efficacy of the adversarial patch in perturbing the model’s spatial perception, inducing a consistent deviation from the intended direction of movement, ultimately resulting in failure in a long run. For TMA, we observe a notable reduction in the range of motion along the x-axis corresponding to the targeted attack axis. This suggests that our proposed targeted attack can effectively constrained robot movements.
In summary, the qualitative analysis shows that both Untarget attack, UADA, UPA, and TMA can effectively disrupt the robot actions generated from VLA models. These findings underscore a pressing security concern during the deployment of generalist robots, especially when considering application scenes that requires reliable operations [11, 68].
Real-world Performance. In addition to the digital simulation results, we also conduct a comprehensive evaluation of the performance of our generated adversarial patches in real-world scenarios. The evaluation encompassed 100 trials across three distinct tasks: object grasping, placement, and manipulation (see §S5 for more details). As shown in Fig. 4, the adversarial patch generated with UADA demonstrated the ability to manipulate the robot effectively, achieving an attack success rate exceeding 43%. Although this success rate is lower than the corresponding digital-world performance (i.e., 100%), it highlights the effectiveness of our patches in physical-world applications as well without the need for further adaptations. Crucially, we observed that the induced erratic movements (similar to simulation scenario) of the robotic manipulators during successful attacks pose significant risks to human safety and surrounding environments. This finding emphasizes the severe threat posed by VLA models in real-world applications, consistent with similar observations in digital settings.
4.4 Diagnostic Experiment
Impact of Inner-loops. We discuss the impact of inner-loop steps to model performance in Fig. 5(a). The results show that NAD first improves when inner-loop steps continue to increase. This is due to the reduced data variance and gradually stabilized gradients. However, further increasing the steps results in a slower performance increase, with a noticeably longer training schedule. We thus choose inner-loop steps=50 as it balances between performance and scale.
Impact of Patch Size. We further study TMA’s performance with different patch sizes and report the average FR across four LIBERO [44] benchmark tasks in Fig. 5. The patch sizes examined were $[3\%,5\%,7\%,10\%]$ of a $224\times 224$ input image. Our findings indicate an inversely proportional relationship between the $L_{1}$ distance of predicted actions and the target action as the patch size increased. This increase in patch size also correlated with a marked rise in FR. The observed trend suggests that larger patch sizes afford the adversary more optimization space to influence the model’s output, aligning with prior work [12, 81].
4.5 Robustness Evaluation
We further examine if concurrent defense strategies [18, 80, 78] can generalize against our adversarial examples introduced within VLA. Specifically, we applied four prior defense techniques and reported their FR for both benign and adversarial samples in Fig. 5, respectively. The results indicate that our adversarial attack can escape from most of the defense strategies. More details see in §S6.
4.6 Systemic Discussion
(a) BridgeData V2 [69]
(b) LIBERO [44]
Figure 6: Patch Visualization. Sematic-rich patches.
Patch Pattern Analysis. In Fig. 6, we present the adversarial patches generated for a range of targets. A particularly noteworthy observation is that some of these patches bear a striking resemblance to the structural joints of a robotic arm. Given the strong similarity between these patches and robotic arms in appearance, we hypothesize that these VLA-based models, in their current training paradigms, are often limited to a single robotic system operating within restricted camera views. This constrained perspective/camera pose may inadvertently induce a learned representation bias, where adversarial perturbations align with prominent structural features of the training robot’s appearance in order to deceive current victim models. Given that these patches are model-specific and reflect the spatial and visual patterns of the training environment, we have a reasonable concern that they may compromise VLA-based models’ generalizability and robustness in physical-world deployments.
Possible Defense Strategies. Our work highlights the pressing need for a paradigm shift in the VLA-based models’ training strategies in order to successfully defend our attacks. Specifically, future approaches might incorporate more training scenarios that involve complex multi-robot interactions to mitigate the current patch bias. Another possible solution is to leverage the logical relationships within the robot arm’s physical structure to estimate the actual position of each joint. By predicting joint positions deviated from physical constraints, we can eliminate unreasonable adversarial samples, even if they are similar to actual joint(s).
5 Conclusion
While VLA models have gained significant popularity due to their substantial robot capabilities, this study, for the first time, establishes specific robotic attack objectives and demonstrates that these models are in fact vulnerable to various types of attacks. demonstrate that our attacks expose significant vulnerabilities to VLA models, raising concerns regarding impetuous real-world deployments. We believe our framework provides pioneering and foundational contributions to the reliability of AI-enhanced robotic systems.
Supplementary Material
SUMMARY OF THE APPENDIX
This appendix contains additional experimental results and discussions of our work, organized as:
  • §S1 includes additional related work in for robotic system safety.
  • §S2 includes additional simulation results for black-box attacks with different VLA-based models.
  • §S3 outlines the experiment details.
  • §S4 provides targeted attack supply results, investigating TMA across various action magnitudes and patch sizes.
  • §S5 explains tasks adopted for real-world experiments.
  • §S6 outlines the parameter configurations used in robustness evaluations.
  • §S7 studies failure cases, providing an in-depth analysis of the limitations of the proposed methods.
  • §S8 supplies additional information on diagnostic experiments.
  • §S9 visualizes more qualitative results.
  • §S10 provides discussions on future directions, highlighting potential areas for further research.
S1 Human-Robot Interaction Safety
Current human-robot interaction risks include physical discomfort or injury [22, 4, 5]. Considering these risks, ensuring safety during interactions becomes the top priority as many literature outlined [67, 24, 39]. Existing robotic safety strategies have primarily concentrated on mitigating naturally occurring hazards [39]. These strategies can be generally categorized into pre-collision or post-collision interventions based on the timing of manage hazards [39]. Pre-collision approaches aim to prevent severe hazards by implementing physical constraints, such as controlling force and speed [28, 30], defining safety zones [68] and predicting human actions to avoid potentially dangerous robot movements [36, 35, 50]. Post-collision strategies, on the other hand, focus on detection and response after a harmful action has occurred [48, 20]. While there is substantial research on robotic safety, the integration of AI into robotic policies poses new challenges, where attackers can manipulate AI-driven robot and execute malicious actions. Such threats introduce additional layers of risk and complexity. In this paper, we primarily focus on adversarial threats.
S2 Black-box Attack Results
This section examines the black-box settings of our proposed attacks, focusing on evaluating the transferability of the proposed patch attack. To this end, we include another VLA-based model in our experiment: LLaRA [42].
Setup. For the LLaRA black-box experiment, we generate adversarial patch from simulation (i.e. openvla-7B-libero-long [34] model with LIBERO Long [44] dataset) and physcial (i.e. openvla-7B [34] model with BridgeData V2 [69] dataset). These patches are then pasted at the top-left corner of the image in the VIMA [32] scenario.
Table S1: LLaRA [42] black-box results. We report failure rates (%) in VIMA [32] scenario. Simulation and physical represent the same experiment setup as provided in §4.2. The FR ($\uparrow$) is highlighted in best and second best for each difficulty.
Task Difficulty Setup Objective DoF L1 L2 L3 L4 LLaRA (D-inBC + Aux (D) + Oracle) 10.0 11.9 20.8 66.2 Simulation Untargeted       DoF1 15.8 20.8 29.2 67.5 Untargeted       DoF1-3 15.8 16.5 27.5 70.0 UADA       DoF1 15.0 18.8 27.1 71.3 UPA       DoF1-3 15.9 19.1 28.8 71.1 Physical Untargeted       DoF1 13.8 23.5 26.2 76.4 Untargeted       DoF1-3 14.2 26.4 27.6 74.7 UADA       DoF1 16.3 27.2 28.5 76.1 UPA       DoF1-3 12.7 25.6 28.7 72.1
Results. Across the four task difficulty tasks (L1-L4) in VIMA [32], adversarial patches significantly increase the failure rate (See Tab. S1) compared to the benign result of LLaRA (D-inBC + Aux (D) + Oracle). For L1, the UPA objective under the simulation setting increases the failure rate by 6.3% (16.3% $v.s.$ 10.0%). For L2, the adversarial patch generated with Untargeted objective under physical setup achieves the highest failure rate, increasing the faliure rate by 15.3% (27.2% $v.s.$ 11.9%). For L3, Untargeted Action Discrepancy Attack (UADA) under the physical setting demonstrates strong effectiveness, increasing the failure rate by 8.4% (29.2% $v.s.$ 20.8%). Finally, for L4, the physical setting with the Untargeted objective effectively increases the failure rate by 10.2% (76.4% $v.s.$ 66.2%).
S3 Experiment Details
Attacking Details: Following [34], we apply scaling and normalization for data preprocessing. The initial learning rate is set to $2e^{-3}$, with the AdamW optimizer [47] with a cosine annealing scheduler, including $20$ warm-up iterations. Training is conducted over $2e^{3}$ iterations with a batch size of $6$ and $50$ inner-loop steps (see Algorithm 1). The transformation parameters $\phi,\psi$ in Eq. 9 are set to $30$ and $0.2$ to simulate real-world perspective change. For all three objectives, considering the trade-off between stealth and performance observed in §S4, we set the patch size to $5\%$. For the hyperparameters $\alpha,\beta$ in Eq. 5, We conduct grid search and tune $\alpha$ from $[0.2,0.4,0.6,0.8]$ and $\beta$ from $[0.8,0.6,0.4,0.2]$. We select $\alpha=0.8$ and $\beta=0.2$ based on the validation set performance. Ablation studies $w.r.t.$ $\alpha$ and $\beta$ are presented in §S8.
Evaluation Details: To rigorously assess the effectiveness of attacks and minimize task failures arising from random placement of adversarial patches that obscure critical task-relevant objects, we identify base left-corner locations for patch placement specific to each of the four tasks in LIBERO [44]. For the Spatial task, the base point is set as $(120,160)$. For the Object task, the base point is $(30,150)$. Similarly, the base point is $(15,158)$ for the Goal task and $(5,160)$ for the Long task. Additionally, to enhance reproducibility by mitigating randomness, patches remain untransformed during evaluation. This approach ensures a reliable assessment of adversarial patches’ impact on task performance.
Figure S1: UADA Qualitative Results The figure illustrates the 3D and 2D trajectories for both benign $\bullet$ and adversarial $\bullet$ scenarios, highlighting the impact of the adversarial patch at each time step. We visualize the start point, marked as , and the endpoint, marked as .
Figure S2: UPA Qualitative Results The figure illustrates the 3D and 2D trajectories for both benign $\bullet$ and adversarial $\bullet$ scenarios, highlighting the impact of the adversarial patch at each time step. We visualize the start point, marked as , and the endpoint, marked as
Figure S3: TMA Qualitative Results The figure illustrates the 3D and 2D trajectories for both benign $\bullet$ and adversarial $\bullet$ scenarios, highlighting the impact of the adversarial patch at each time step. We visualize the start point, marked as , and the endpoint, marked as
S4 Targeted Attack supply Results
To further explore the action manipulation capabilities of TMA, we conduct experiments targeting varying magnitudes to assess its effectiveness across different scenarios.
Setup. We conduct experiments attack at DoF1 and utilize various action magnitudes (i.e. $0.5,-0.5,1.0,-1.0$) as the attack targets under simulation and physical settings.
Table S2: Different Magnitude results. Average failure rate (%) is reported in 4 LIBERO simulation tasks. The initial AFR across the four tasks is 23.5%.
Action Magnitude Target Action Metric 0.5 -0.5 1.0 -1.0 Simulation L1 0.072 0.049 0.132 0.107 AFR 82.5 87.8 78.4 74.2 Physical L1 0.055 0.071 0.103 0.121 AFR 62.5 68.8 63.8 58.3
Results. The results are presented in Tab. S2. In the simulation setup, the highest AFR (87.8%) is achieved at a target magnitude of $-0.5$, representing a substantial increase of 64.3% compared to the initial AFR (87.8% $v.s.$ 23.5%). Similarly, in the physical setup, the target magnitude of $-0.5$ yields the highest AFR of 68.8%. These findings highlight the superior manipulation capability and disruptive effectiveness of TMA.
Figure S4: Redundant DoF in a Failure Case. The first row is the benign scenario, while the second row illustrates the adversarial scenario. In the adversarial scenario, the adversarial patch is generated by targeting DoF4 within the simulation setup. In this task, orientation (i.e. DoF4-6) is identified as a redundant DoF since the task completion does not require any changes in orientation. As a result, the DoF4 attack, which focuses on orientation, fails to disrupt the task’s execution.
S5 Details of Real-world Experiments
We design four tasks to evaluate the impact of adversarial samples on task failure rates in real-world scenarios. The specific tasks include “put the carrot on the plate”, “put the corn on the plate”, “put the carrot into the bowl”, and “flip the pot upright”. To fairly evaluate the performance of our attack, we only consider tasks where the VLA model operates successfully, excluding failure cases. We then introduce the adversarial patch to assess its impact. This leads to an average task failure rate of 43%, underscoring the significant disruptive potential of adversarial samples on robotic tasks.
S6 More Robustness Evaluation Details
In this section, we provide a detailed parameter setting of conducted robustness evaluation in §4.5. Specifically, we assess the effectiveness of four commonly employed defense techniques, including JPEG compression [18], bit-depth reduction [78], median blur [78], and Gaussian noise [80]. JPEG compression applies compression algorithms to the input images prior to feeding them into the depth estimation network, aiming to disrupt adversarial patterns. We test compression quality levels ranging from 50 to 10, with lower quality levels corresponding to higher compression rates. Gaussian noise introduces zero-mean Gaussian noise to the input image, leveraging its randomness to counteract the structured nature of adversarial perturbations. The standard deviation of the noise varies from 0.01 to 0.1, with higher values introducing stronger noise. Median blur smooths the image by replacing each pixel value with the median of its surrounding pixels, using square kernel sizes from 3 to 9; larger kernel sizes produce stronger smoothing effects. Bit-depth reduction remaps the standard 8-bit depth of RGB channels to smaller bit depths, reducing the color space and potentially disrupting adversarial perturbations. We evaluate cases with bit depths ranging from 6 bits to 3 bits. These defense techniques are evaluated to understand the robustness of our attack (results see in Fig. 5).
S7 Failure Case Analysis
This section analyzes the attack failure cases to gain deeper insights into robot manipulation and highlight the critical role of DoF targets on task failure.
We analyze most of attack failure cases and find that certain DoFs are redundant in task execution. In Fig. S4, the adversarial patch targeting DoF4 with a value of 0 fails to disrupt the execution of the task. This failure can be attributed to the fact that DoF4 controls the orientation along the x-axis, which is redundant for grasping objects such as bowls in the context of this task. As a result, attacks targeting redundant DoF of the task [38] are less likely to disrupt successful execution effectively. This observation underscores the importance of task-specific considerations when designing adversarial attacks on robotic systems.
Table S3: Parameter Diagnostic Results. Four distinct $\alpha$ and $\beta$ combination results. We report the average failure rate of four tasks in LIBERO [44].
Setup $\alpha$ $\beta$ AFR (%) Simulation 0.2 0.8 86.5 0.4 0.6 90.5 0.6 0.4 89.4 0.8 0.2 93.4
S8 More Diagnostic Experiment Details
We perform diagnostic experiments to assess the influence of the hyperparameters $\alpha$ and $\beta$ in UPA. As shown in Tab. S3, when $\alpha=0.8$ and $\beta=0.2$, UPA achieves the highest average score of 93.4 across the four LIBERO [44] tasks among all tested configurations. This result indicates that this specific combination optimally balances the trade-offs between direction and magnitude adjustments governed by $\alpha$ and $\beta$, establishing it as the most effective choice.
S9 More Qualitative Results
This section presents additional qualitative results for each of the three attack objectives, complementing results in §4.3 and offering deeper insights into the effectiveness of our adversarial attacks. Specifically, the results of UADA, UPA, and TMA are shown in Fig. S1S2, and S3, respectively. These qualitative results reveal significant deviations in the adversarial trajectories of the proposed methods across all four LIBERO tasks compared to the benign trajectories. Specifically, UADA induces substantial action discrepancies while diverging sharply from the benign trajectory, highlighting its disruptive impact. For UPA, a significant position deviation is observed on the spatial task [44], where the adversarial trajectory deviates significantly from the benign trajectory, ultimately causing task failure. In contrast, the TMA exhibits a smaller action amplitude, successfully manipulating the robot’s behavior and resulting in task failure. These findings highlight the distinct impacts and underlying mechanisms of the three attack methods.
S10 Future direction
Future research on attacks against VLA-based models can potentially be conducted on two key objectives: enhancing camouflage and ensuring practical feasibility. Improving camouflage involves reducing detection probability by generating adversarial patches that seamlessly integrate with the environment, leveraging natural patterns and context-aware design. Furthermore, future efforts should avoid targeting redundant DoF and instead concentrate on critical components that are most likely to disrupt task performance. This can be achieved by leveraging task-specific knowledge and employing advanced optimization techniques to maximize the effectiveness of such attacks while aligning with real-world physical constraints.